# Prym loci of branched double coverings and generalized Andreotti-Mayer loci

### v1

@@ 4. Andreotti-Mayer loci
is an analytic subset in $\mathcal{S}$, and by the proper mapping theorem ([6, 10.6.1.]), $\mathcal{N}_{m}=\upsilon(\mathcal{S}_{m})$ is an analytic subset in $\mathfrak{H}_{d}$. Let $\mathcal{N}$ be a local irreducible component of $\mathcal{N}_{m}$ at $\tau_{0}\in\mathcal{N}_{m}$. We denote by $\Lambda(\mathcal{N})$ the set of all irreducible components $\mathcal{M}$ of $\upsilon^{-1}(\mathcal{N})\cap\mathcal{S}_{m}$ such that $\upsilon(\mathcal{M})=\mathcal{N}$. Since $\upsilon$ is proper, $\Lambda(\mathcal{N})$ is a non-empty finite set. Following [4], we set $\mathcal{S}(\tau_{0},\mathcal{N})=\upsilon^{-1}(\tau_{0})\cap\bigcup_{\mathcal{M}\in\Lambda(\mathcal{N})}\mathcal{M}$.
@@
$T_{\mathcal{N},\tau_{0}}\subset\left(\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{a}},\tau_{0})\,d\tau_{kl}\right)^{\perp}\subset{T_{\mathfrak{H}_{d},\tau_{0}}}$
@@
is a non-empty open subset of $\mathcal{M}$. For $({\boldsymbol{z}},\tau)\in{U}$, we can take a local coordinate ${\boldsymbol{t}}=(t_{1},\dots,t_{\dim{\mathcal{N}}})$ of $\mathcal{N}$ at $\tau$ and a local coordinate $({\boldsymbol{s}},{\boldsymbol{t}})$ of $U$ at $({\boldsymbol{z}},\tau)$ such that $\upsilon:U\rightarrow\mathcal{N}$ is given by $({\boldsymbol{s}},{\boldsymbol{t}})\mapsto{\boldsymbol{t}}$. Let $({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\in{U}\subset\mathcal{S}$ be the point corresponding to $({\boldsymbol{s}},{\boldsymbol{t}})$. Then we have
[…]
@@
=$\mathcal{M}^{\prime}=\left\{({\boldsymbol{z}},\tau)\in\mathcal{M}\,\middle|\,T_{\mathcal{N},\tau}\subset\Bigl{(}\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{z}},\tau)\,d\tau_{kl}\Bigr{)}^{\perp}\right\},$
@@
$\sum_{1\leq{k}\leq{l}\leq{d}}\sum_{{\boldsymbol{i}}\in{I}}c_{{\boldsymbol{i}},j}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{a}}_{j},\tau_{0})\,d\tau_{kl}\bigoplus_{1\leq{k}\leq{l}\leq{d}}\mathbb{C}d\tau_{kl}\quad(1\leq{j}\leq{r})$
@@
Let $\pi:\mathfrak{H}_{d}\rightarrow\mathcal{A}_{d}^{\Delta}=\Gamma_{\Delta}\backslash\mathfrak{H}_{d}$ be the quotient map. By Corollary 2.8, we have $\pi^{-1}(\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})})\subset\mathcal{N}_{n-4}$ for $n\geq 4$, and $\dim{\upsilon^{-1}(\tau_{0})}=n-4$ for any $\tau_{0}\in\pi^{-1}(\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})})$.
@@
When $n\geq 5$, by Corollary 2.8, $\upsilon^{-1}(\tau_{0})$ is irreducible of dimension $n-4$. Since
$n-4=\dim{\upsilon^{-1}(\tau_{0})}\geq\dim_{({\boldsymbol{z}},\tau_{0})}{(\upsilon^{-1}(\tau_{0})\cap\mathcal{M})}\geq{n-4},$
for any irreducible component $\mathcal{M}$ of $\upsilon^{-1}(\mathcal{N})\cap\mathcal{S}_{m}$ and $({\boldsymbol{z}},\tau_{0})\in\mathcal{M}$. We have $\upsilon^{-1}(\tau_{0})\subset\mathcal{M}$ for any $\mathcal{M}\in\Lambda(\mathcal{N})$, hence $\upsilon^{-1}(\tau_{0})\subset\mathcal{S}(\tau_{0},\mathcal{N})$. When $n=4$, by Lemma 4.1, for $\mathcal{M}\in\Lambda(\mathcal{N})$ and ${\boldsymbol{i}}\in{I}$, there exists $\mathcal{M}_{\boldsymbol{i}}\in\Lambda(\mathcal{N})$ such that
$\mathcal{M}_{\boldsymbol{i}}=\{({\boldsymbol{z}},\tau)\in\mathbb{C}^{d}\times\mathfrak{H}_{d}\mid({\boldsymbol{z}}-\tau\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix},\tau)\in\mathcal{M}\}.$
@@
$\mathcal{R}_{g,2n}^{h}=\{[\tilde{C}\overset{\phi}{\rightarrow}C]\in\mathcal{R}_{g,2h}\mid\text{$C$ is a hyperelliptic curve}\}$
@@
$\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{1}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{j})z_{k}z_{l},\dots,\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{r}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{j})z_{k}z_{l}\in\mathbb{C}[z_{1},\dots,z_{d}].$
@@
$\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{1}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{j})d\tau_{kl},\dots,\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{r}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{j})d\tau_{kl}\in\bigoplus_{1\leq{k}\leq{l}\leq{d}}\mathbb{C}d\tau_{kl}$
@@
Since $\mathcal{R}_{g,2n}$ is irreducible by [3, p138], there exist an irreducible component $\mathcal{P}\subset\mathcal{A}_{d}^{\Delta}$ of the generalized Andreotti-Mayer locus $\mathcal{N}_{d,n-4}^{\Delta}$ which contains $\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})}$. Then, we have $\pi^{-1}(\mathcal{P})\subset\mathcal{N}_{n-4}$, and

### v2

@@ 4. Andreotti-Mayer loci
is an analytic subset in $\mathcal{S}$, and by the proper mapping theorem ([6, 10.6.1.]), $\mathcal{N}_{m}=\upsilon(\mathcal{S}_{m})$ is an analytic subset in $\mathfrak{H}_{d}$. Let $\mathcal{N}$ be a local irreducible component of $\mathcal{N}_{m}$ at $\tau_{0}\in\mathcal{N}_{m}$. We denote by $\Lambda(\mathcal{N})$ the set of all irreducible components $\mathcal{M}$ of $\upsilon^{-1}(\mathcal{N})\cap\mathcal{S}_{m}$ such that $\upsilon(\mathcal{M})=\mathcal{N}$. Since $\upsilon$ is proper, $\Lambda(\mathcal{N})$ is a non-empty finite set. Following [4], we consider the subset $\mathcal{S}(\tau_{0},\mathcal{N})=\upsilon^{-1}(\tau_{0})\cap\bigcup_{\mathcal{M}\in\Lambda(\mathcal{N})}\mathcal{M}$ in $\upsilon^{-1}(\tau_{0})$.
@@
$T_{\mathcal{N},\tau_{0}}\subset\Bigl(\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{a}},\tau_{0})\,d\tau_{kl}\Bigr)^{\perp}\subset{T_{\mathfrak{H}_{d},\tau_{0}}}$
@@
is a non-empty open subset of $\mathcal{M}$. For $({\boldsymbol{z}},\tau)\in{U}$, we can take a local coordinate ${\boldsymbol{t}}=(t_{1},\dots,t_{\dim{\mathcal{N}}})$ of $\mathcal{N}$ at $\tau$ and a local coordinate $({\boldsymbol{s}},{\boldsymbol{t}})$ of $U$ at $({\boldsymbol{z}},\tau)$ such that $\upsilon|_{U}:U\rightarrow\mathcal{N}$ is given by $({\boldsymbol{s}},{\boldsymbol{t}})\mapsto{\boldsymbol{t}}$. Let $({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\in{U}\subset\mathcal{S}$ be the point corresponding to $({\boldsymbol{s}},{\boldsymbol{t}})$. Then we have
$\theta_{\boldsymbol{i}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))=0,\quad\frac{\partial\theta_{\boldsymbol{i}}}{\partial{z_{j}}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))=0\quad({\boldsymbol{i}}\in{I},\,1\leq{j}\leq{d}).$
@@
$\mathcal{M}^{\prime}=\left\{({\boldsymbol{z}},\tau)\in\mathcal{M}\,\middle|\,T_{\mathcal{N},\tau}\subset\Bigl(\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{z}},\tau)\,d\tau_{kl}\Bigr)^{\perp}\right\},$
@@
$\sum_{1\leq{k}\leq{l}\leq{d}}\sum_{{\boldsymbol{i}}\in{I}}c_{{\boldsymbol{i}},j}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{a}}_{j},\tau_{0})\,d\tau_{kl}\in{T_{\mathfrak{H}_{d},\tau_{0}}^{\vee}}=\bigoplus_{1\leq{k}\leq{l}\leq{d}}\mathbb{C}d\tau_{kl}\quad(1\leq{j}\leq{r})$
@@
[…]
@@
When $n\geq 5$, by Corollary 2.8, $\upsilon^{-1}(\tau_{0})$ is irreducible of dimension n4n-4. Since
$n-4=\dim{\upsilon^{-1}(\tau_{0})}\geq\dim_{({\boldsymbol{z}},\tau_{0})}{(\upsilon^{-1}(\tau_{0})\cap\mathcal{M})}\geq{n-4}$
for any irreducible component $\mathcal{M}$ of $\upsilon^{-1}(\mathcal{N})\cap\mathcal{S}_{m}$ and $({\boldsymbol{z}},\tau_{0})\in\mathcal{M}$, we have $\upsilon^{-1}(\tau_{0})\subset\mathcal{M}$ for any $\mathcal{M}\in\Lambda(\mathcal{N})$, hence $\upsilon^{-1}(\tau_{0})\subset\mathcal{S}(\tau_{0},\mathcal{N})$. When $n=4$, by Lemma 4.1, for $\mathcal{M}\in\Lambda(\mathcal{N})$ and ${\boldsymbol{i}},{\boldsymbol{j}}\in{I}$, there exists $\mathcal{M}_{{\boldsymbol{i}},{\boldsymbol{j}}}\in\Lambda(\mathcal{N})$ such that
$\mathcal{M}_{{\boldsymbol{i}},{\boldsymbol{j}}}=\bigl\{({\boldsymbol{z}},\tau)\in\mathbb{C}^{d}\times\mathfrak{H}_{d}\mid({\boldsymbol{z}}-\tau\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix}-\Delta\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{j}}\end{pmatrix},\tau)\in\mathcal{M}\bigr\}.$
@@
$\mathcal{R}_{g,2n}^{h}=\{[\tilde{C}\overset{\phi}{\rightarrow}C]\in\mathcal{R}_{g,2h}\mid\text{$C$ is a hyperelliptic curve}\}.$
@@
$\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{1}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{1})z_{k}z_{l},\dots,\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{r}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{r})z_{k}z_{l}\in\mathbb{C}[z_{1},\dots,z_{d}].$
@@
$\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{1}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{1})d\tau_{kl},\dots,\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{r}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{r})d\tau_{kl}\in\bigoplus_{1\leq{k}\leq{l}\leq{d}}\mathbb{C}d\tau_{kl}$
@@
Since $\mathcal{R}_{g,2n}$ is irreducible by [3, p138], there exist an irreducible component $\mathcal{P}$ of the generalized Andreotti-Mayer locus $\mathcal{N}_{d,n-4}^{\Delta}\subset\mathcal{A}_{d}^{\Delta}$ such that $\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})}\subset\mathcal{P}$. Then, we have $\pi^{-1}(\mathcal{P})\subset\mathcal{N}_{n-4}$, and